# Neural ordinary differential equations for model order reduction of stiff systems

### v1

@@ 2.3 Data-driven model order reduction
Figure 1: Example of discrete solution manifold $\mathcal{M}_{h}$ for the Van der Pol oscillator.
@@
=$\mathcal{M}=\{\boldsymbol{u}(t;\boldsymbol{\mu})\,|\,t\in[0,T),\boldsymbol{\mu}\in\Gamma\}\subset\mathbb{R}^{N_{u}},$ (3)
@@
which is a collection of the solutions for each parameter $\boldsymbol{\mu}\in\Gamma$ and time $t\in[0,T)$. We suppose that it is possible to find a discretized approximation $\boldsymbol{u}_{h}(t;\boldsymbol{\mu})$ up to arbitrary accuracy by integrating (1) with a suitable method. By solving the problem for $M_{\mu}$ different choices of $\boldsymbol{\mu}$ we create a discretized manifold
$\mathcal{M}_{h}=\{\boldsymbol{u}_{h}(t;\boldsymbol{\mu})\,|\,t\in\{0\leq t_{1}<...<t_{M_{T}}\leq T\},\boldsymbol{\mu}\in\Gamma_{h}=\{\boldsymbol{\mu}_{1}\in\Gamma,...,\boldsymbol{\mu}_{M_{\mu}}\in\Gamma\}\}\subset\mathbb{R}^{N_{u}},$ (4)
@@
which is a finite set of vectors $\boldsymbol{u}_{h}\in\mathbb{R}^{N_{u}}$ that approximate the solution at a certain discretized point in time $t$ and for a certain choice of parameter $\boldsymbol{\mu}$. This set can be reshaped into a so-called, snapshot matrix $\boldsymbol{U}\in\mathbb{R}^{N_{u}\times M}$, where $M=|\mathcal{M}_{h}|$, by stacking all the column vectors of the discrete manifold. Let us stress that the time discretization $\{0\leq t_{1}<...<t_{M_{T}}\leq T\}$ may change depending on the value of $\boldsymbol{\mu}$. Indeed, it often happens that $\boldsymbol{\mu}$ changes the dynamics of the system, and thus a different time discretization is generated by the adaptive procedure. For the sake of simplicity, we omit the explicit dependence on $\boldsymbol{\mu}$ of the time discretization.
In practice, one might obtain data also from knowing the exact analytical solution of the problem for some specific values of $\boldsymbol{\mu}$ or from real-world measurements. However, it is of utmost importance that the data gathered in the snapshot matrix $\boldsymbol{U}$ is reliable since it is interpreted as our “ground truth”.
@@
Data-driven ROMs aim to leverage the information contained in the discretized solution manifold $\mathcal{M}_{h}$ to obtain in a fast and reliable way the solution $\boldsymbol{u}_{h}(t;\mu)$ for $\boldsymbol{\mu}\notin\Gamma_{h}$ or $t\geq T$. Figure 1 shows a graphical representation of the solution manifold for the Van der Pol oscillator (cf. Section 4.1). The motivation behind ROMs is that classical time integration methods may become prohibitively expensive either due to the size of $N_{u}$ or due to stiffness (cf. Section 3).
2.4 Neural networks
In this section, we review the basic concepts of deep learning that we will use to build our ROM. Let $N_{I},N_{O}\in\mathbb{N}$ be strictly positive. An artificial neural network is a function $\mathcal{F}\,:\,\mathbb{R}^{N_{I}}\rightarrow\mathbb{R}^{N_{O}}$ that maps an input vector $\boldsymbol{x}\in\mathbb{R}^{N_{I}}$ to an output vector $\boldsymbol{y}\in\mathbb{R}^{N_{O}}$ and depends on the set of parameters ${\boldsymbol{\theta}}$ (we think of all the parameters as reshaped into a one-dimensional vector). As architecture, we consider the dense fully connected feedforward (FNN) neural network (also called multi-layer perceptron) which is defined as follows. Fix two positive integers $L$ and $Z$. Let $\boldsymbol{y}^{(0)}=\boldsymbol{x}$ and $\boldsymbol{y}^{(L)}=\tilde{\boldsymbol{y}}$. A FNN of depth $L$ and width $Z$ is the composition of $L$ functions called layers defined by
$\boldsymbol{y}^{(l)}=\sigma^{(l)}(\mathrm{W}^{(l)}\boldsymbol{y}^{(l-1)}+\boldsymbol{b}^{(l)}),\quad l=1,...,L,$
where $\mathrm{W}^{(l)}\in\mathbb{R}^{N_{l}\times N_{l-1}}$ are matrices of parameters called weights and $\boldsymbol{b}^{(l)}\in\mathbb{R}^{N_{l}}$ are vectors of parameters called biases. Here, $N_{0}=N_{I}$, $N_{L}=N_{O}$ and $N_{l}=Z$ for $l=1,...,L-1$. Finally, at the end of each layer, a scalar non-linear activation function $\sigma^{(l)}$ is applied component-wise. We employ the same activation function (to be specified later, cf. Section 4) for all the layers apart from the last one where we use a linear activation function.
Several learning strategies are available to determine the vector of parameters $\boldsymbol{\theta}$. Among these, one of the most popular paradigms is supervised learning. This method focuses on defining and minimizing a non-linear cost function, known as the loss function $\mathcal{L}$. The loss function is typically formulated using a dataset comprising input-target pairs ${(\boldsymbol{x},{\boldsymbol{y}})}_{i=1}^{N{\textnormal{train}}}$, which the neural network aims to learn. A common expression for the loss function is:
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{i=1}^{N_{\textnormal{train}}}\left\|{\boldsymbol{y}}-\tilde{\boldsymbol{y}}(\boldsymbol{x};\boldsymbol{\theta})\right\|_{p},$ (5)
where $\left\|\cdot\right\|_{p}$ denotes the discrete $p$-norm, typically with $p=1$ or $p=2$.
To minimize the loss function gradient-based optimization is commonly employed. The gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}$ is efficiently computed using automatic differentiation techniques [51]. An epoch, defined as one complete pass through the dataset, consists of multiple gradient updates and is the unit often used to quantity the training length. Properly optimized neural networks exhibit the ability to generalize, making accurate predictions on unseen data. For further reading on neural networks, comprehensive resources include [27, 71].
@@ 2.5 Neural ODEs
The model that we choose to use to build the ROM is neural ODEs. Neural ODEs have emerged as powerful tools to model continuous-time dynamics using a differential equation. This approach provides several key benefits. Firstly, Neural ODEs offer superior flexibility in modeling time-series data, as they can naturally handle irregularly sampled data and varying time steps. Secondly, they ensure smooth trajectories and better integration with physical laws and other domains where the underlying processes are continuous in nature. Furthermore, neural ODEs can lead to more efficient training and inference, as they require fewer parameters and can leverage advanced numerical solvers for differential equations. Neural ODEs are also more memory efficient compared to traditional recurrent neural networks, especially for long time-series data, because they avoid the need to store intermediate states of the network during backpropagation.
More precisely, a neural ODE is a neural network NN that evolves the dynamics as
$\dot{\boldsymbol{u}}(t)=\mathcal{NN}(t,\boldsymbol{u}(t);\boldsymbol{\mu}),$ (6)
that is the NN is employed to approximate the right-hand side of the dynamical system (1) by taking as inputs the time $t$, the current state $\boldsymbol{u}$, and the parameters $\boldsymbol{\mu}$. One of the main drawbacks of this approach is that it is necessary to numerically integrate the system in the online phase. Given the recurrent nature of the architecture, adjoint solvers have been developed to enhance the memory efficiency during the training phase [19]; their cost however remains much larger when compared to FNN.
When used as an architecture for ROMs, the neural ODE is often paired with another neural network to reduce the dimension of the system from $N_{u}$ to a small latent state space. This is usually achieved by means of autoencoders [50, 65, 9, 47, 26] or other reduction techniques [59]. However, if $N_{u}$ is small the advantages of using a small latent space representation may not be counterbalanced by the increased cost of evaluating the neural network. In this work, we propose a solution to this latter problem by introducing an appropriate change of variable that significantly reduces the stiffness of the problem. Indeed, we will show that our method makes it possible to use explicit solvers to integrate neural ODEs representing stiff systems.
We report in Figure 2 a representation of the neural ODE. The hat symbol $\hat{\cdot}$ represents a suitable normalization of the data, see the next section for details. For the sake of simplicity, we will indicate in the captions the reference solution (the one that is comprised in the dataset) with the subscript “$ref$”, and the components of $\boldsymbol{u}$ as predicted by the neural ODE with simply $u_{i}$.
@@ 2.5.1 Training a Neural ODE
There are two main approaches to training a neural ODE. The first one is the classic supervised approach. Assuming that the value of $\boldsymbol{f}(t,\boldsymbol{u}_{h}(t);\boldsymbol{\mu})$ is known for each sample in the discrete manifold $\boldsymbol{u}_{h}\in\mathcal{M}_{h}$, then it is possible to compute the loss (5) and minimize it through a gradient-based optimizer. However, this kind of data is not always available. Instead, the second approach – the one originally proposed for the training of neural ODEs – minimizes the error with respect to the timeseries data $\boldsymbol{u}_{h}(t)$. Namely, by using a differentiable ODE solver, we minimize the loss
@@
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{i=1}^{N_{\textnormal{train}}}\left\|{\boldsymbol{u}_{h}(t;\boldsymbol{\mu})}-\left(\boldsymbol{u}_{h}(0;\boldsymbol{\mu})+\int_{0}^{t}\mathcal{NN}(\boldsymbol{u}(s;\boldsymbol{\mu}),\boldsymbol{\mu})\mathrm{d}s\right)\right\|_{p},$ (7)
@@
where the integral is usually computed with an explicit integration method like forward Euler or an explicit Runge-Kutta method.
@@
The advantage of the supervised approach is that it is less computationally demanding since it does not need to compute the gradient through the ODE solver using automatic differentiation. On the other hand, in the supervised framework, the loss is not always indicative of the accuracy of the predictions on the timeseries data $\boldsymbol{u}(t)$. For this reason, we always employ as validation loss Eq. (7).
@@
Implementation details for Eq. (7) are important when it is employed to train the neural network. We resample the discrete timeseries $\boldsymbol{u}_{h}(t;\boldsymbol{u})$ to match the output of the time integrator, namely a Runge-Kutta method of the fourth order with a fixed timestep. To minimize the interpolation errors, we use a reference solution with a very fine time discretization. A uniform resampling allows for an efficient usage of the partial results of the ODE solver for all the times smaller than the final one. Moreover, the initial integration time is not zero but is picked uniformly between $[0,T)$ so that each time integration in the dataset is done with the same number of timesteps: this procedure enables vectorization of the computations, greatly reducing the computational cost. This optimization was originally introduced in torchdiffeq [19]. The number of timesteps used is called unroll length and is a hyperparameter that plays a key role in the optimization procedure. Finally, as commonly done in supervised learning, the computations are split into mini-batches to increase stochasticity. If not otherwise stated, when using Eq. (7) for validation or testing we employ as a time integrator an embedded Runge-Kutta method of order four with adaptive time-stepping. However, computations are significantly cheaper since we can turn off automatic differentiation.
@@
In our experiments, we employ a two-phase training. First, a standard supervised approach is used to “initialize” the neural network since it is cheap and enables a fast reduction of the loss. Then, we switch to the differentiable ODE solver approach to fine-tune the ROM. Our experiments show that in this case, it is beneficial to use an increasingly larger unroll length to ensure the robustness and accuracy of the model. Further details will be given in Section 4 for each test case. Employing a two-phase training approach not only reduces the computational cost but also increases the quality of the final model. Indeed, from our experiments, we have observed that with the supervised learning approach the neural network is able to learn sharper gradients that may lead to unstable states. On the other hand, learning through the ODE solver produces stable results that might follow less precisely the reference solution. Figure 3 shows an example for two (underfitting) models. In Figure 4 we show the training and validation loss for a model trained in sequence with the supervised and ODE solver approaches.
@@
To accelerate the training suitable normalizations are applied to the input $\boldsymbol{\mu}$, the output $\boldsymbol{f}$, and the timeseries data $\boldsymbol{u}$. The exact expressions for the normalizations depend on the test case considered and are detailed in Section 4. The insight behind the choice of the normalization is that we aim to have data that has a magnitude close to one, does not have outliers, and has either a uniform distribution or a normal distribution with a standard deviation of magnitude close to one. In our experiments, strictly monotone normalizations achieved better results. We designed the normalizations by manual inspection of the histograms of the data.
@@ 3 Addressing the stiffness issue
While many techniques have been developed to build reduced-order models of large systems stemming from semi-discretized PDEs, not many solutions are available for stiff ODEs. Indeed, it is a challenge to learn their complex dynamics, which often feature large gradients and widely different time scales. More rigorously, a commonly used formula to define the stiffness index of a problem is
$S=\frac{Re(\lambda_{\max{}})}{Re(\lambda_{\min{}})}T,$ (8)
@@
where $\lambda_{i}$ are the eigenvalues of the Jacobian of the system. However, it is known that this definition is not always helpful. For instance, it does not apply to the classic Robertson chemical kinetics problem since it has two zeros eigenvalues [61]. A more pragmatic definition is given by Harier: “Stiff equations are problems for which explicit methods do not work[69]. That is, explicit numerical time integrators such as the embedded Runge-Kutta method of the fourth order require timesteps so small that implicit methods become cheaper. Stiffness is a particularly important issue when working with neural ODEs. Indeed, it is a well-known fact that neural ODE training may fail due to stiffness. For instance, [38] shows issues of this kind and proposes techniques to improve the training. However, it is still necessary to use a costly implicit solver in the online phase, since the learnt dynamic is stiff.
@@
Hence, to use the neural ODE as a ROM, we aim to reduce the stiffness of the system, allowing for the usage of (cheap) explicit methods. To this end, we propose to change the dynamics of the problem by considering the change of variable in time induced by the adaptive time-stepping of the implicit solver used to solve the full-order model. We define a data-driven reparametrization of the time $t$ called $t_{s}=t_{s}(t)$ such that
$t_{s}(t^{n})=\frac{n}{NT}.$
@@
That is, the $\Delta t$ chosen at the $n$-th step by the implicit time stepper is proportional to the derivative of the function that we employ as a change of variable in time. We remark that $t_{s}=t_{s}(t)$ is strictly increasing and thus invertible. We guess this change of variables induces less stiff dynamics since the adaptive time-stepping spreads the sharp peaks and quickly adapts to the large timescales. The system becomes
$\begin{cases}\dot{\boldsymbol{u}}(t_{s})=\boldsymbol{f}_{s}(t_{s},\boldsymbol{u};\boldsymbol{\mu}),\qquad t\in(0,1)\\ \boldsymbol{u}(0;\boldsymbol{\mu})=\boldsymbol{u}_{0}(\boldsymbol{\mu}).\end{cases}$
@@
To accommodate this change in the neural ODE framework, we need a way to compute $t=t(t_{s};\boldsymbol{\mu})$ in the online phase, to map the solution back to the original dynamics. Our experiments show that learning directly the map $t=t(t_{s};\boldsymbol{\mu})$ is a hard task for a neural network, probably due to the presence of abrupt changes in the function. Instead, we found that learning $\dot{t}(\boldsymbol{u}(t_{s};\boldsymbol{\mu}))$ is a stable approach that produces more accurate results. We also prefer this approach to learning $\dot{t}(t_{s};\boldsymbol{\mu})$ to keep the system autonomous (that is, never introduce explicit dependence on time) to improve the generalization properties. The evaluation of the time dynamics is cheap since it can be done in one batch using vectorization. Indeed, in the first step, we compute $\boldsymbol{u}(t_{s})$, then, we carry out the cumulative integration of $\dot{t}(\boldsymbol{u}(t_{s};\boldsymbol{\mu}))$ with the Simpson’s rule. The main advantage of this approach is that it is cheap compared to autoencoders, especially for small problems since the only overhead is the computation of $t=t(t_{s};\boldsymbol{\mu})$. However, it is also possible to combine the autoencoder with the described methodology: first, the autoencoder extracts a latent state representation, then, we could apply the time reparametrization in the latent space, reducing the stiffness of the latent dynamics.

### v2

@@ 2.3 Data-driven model order reduction
Figure 1: Example of discrete solution manifold $\widetilde{\mathcal{M}}$ for the Van der Pol oscillator.
@@
=$\mathcal{M}=\{\boldsymbol{u}(t;\boldsymbol{\mu})\,|\,t\in[0,T],\boldsymbol{\mu}\in\Gamma\}\subset\mathbb{R}^{N_{u}},$ (3)
@@
which is a collection of the values of the solutions for each parameter $\boldsymbol{\mu}\in\Gamma$ and time $t\in[0,T]$. We suppose that it is possible to find a discretized approximation $\boldsymbol{u}(t^{(n)};\boldsymbol{\mu})$ for some times $t^{(n)}\in[0,T]$ up to arbitrary accuracy by integrating (1) with a suitable method. By solving the problem for $N_{\mu}$ different choices of $\boldsymbol{\mu}$ we create a discretized manifold
[…]
@@
which is a finite set of vectors that approximate the solution at a certain discretized point in time $t^{(n)}$ and for a certain choice of parameter $\boldsymbol{\mu}^{(j)}$. Let us stress that the time discretization $\{0\leq t_{1}<...<t_{N(\boldsymbol{\mu}^{(j)})}\leq T\}$ may change depending on the value of $\boldsymbol{\mu}^{(j)}$, and need not be uniform. Indeed, it often happens that $\boldsymbol{\mu}^{(j)}$ changes the dynamics of the system, and thus a different time discretization is generated by the adaptive procedure. For the sake of simplicity, we omit the explicit dependence on $\boldsymbol{\mu}^{(j)}$ of the time discretization.
@@
In practice, one might obtain data also from knowing the exact analytical solution of the problem for some specific values of $\boldsymbol{\mu}$ and $t$ or from real-world measurements. Data-driven ROMs aim to leverage the information contained in the discretized solution manifold $\widetilde{\mathcal{M}}$ to obtain in a fast and reliable way the solution $\boldsymbol{u}(t^{(n)};\boldsymbol{\mu}^{(j)})$ for $\boldsymbol{\mu}^{(j)}$ or $t^{(n)}$ not in the discrete manifold used for training the model. Figure 1 shows a graphical representation of the solution manifold for the Van der Pol oscillator (cf. Section 4.1). The motivation behind ROMs is that classical time integration methods may become prohibitively expensive either due to the size of $N_{u}$ or due to stiffness (cf. Section 3).
@@ 2.4 Neural ODEs
The model that we choose to use to build the ROM is neural ODEs [20]. Neural ODEs have emerged as powerful tools to model continuous-time dynamics using a differential equation. This approach provides several key benefits. Firstly, neural ODEs offer superior flexibility in modeling time-series data, as they can naturally handle irregularly sampled data and varying time steps. Secondly, they ensure smooth trajectories and better integration with physical laws and other domains where the underlying processes are continuous in nature. Furthermore, neural ODEs can lead to more efficient training and inference, as they require fewer parameters and can leverage advanced numerical solvers for differential equations. Neural ODEs are also more memory efficient compared to traditional recurrent neural networks, especially for long time-series data, because they avoid the need to store intermediate states of the network during backpropagation.
More precisely, a neural ODE is a dense fully connected feedforward (FNN) neural network $\mathcal{NN}$ [29, 81], with parameters $\boldsymbol{\theta}$, that evolves the dynamics as
$\dot{\boldsymbol{u}}(t;\boldsymbol{\mu})=\mathcal{NN}(t,\boldsymbol{u}(t);\boldsymbol{\mu}),$ (5)
that is, the FNN is employed to approximate the right-hand side $\boldsymbol{f}$ of the dynamical system (1), thus it takes as inputs the time $t$, the current state $\boldsymbol{u}$, and the parameters $\boldsymbol{\mu}$. In other words, this technique aims to predict $\boldsymbol{u}$ by using $\mathcal{NN}$ to learn the $\boldsymbol{f}$ of Eq. (1). One of the main drawbacks of this approach is that it is necessary to numerically integrate the system in the online phase. Given the recurrent nature of the architecture, adjoint solvers have been developed to enhance the memory efficiency during the training phase [20]; their cost however remains much larger when compared to FNN.
When used as an architecture for ROMs, the neural ODE is often paired with another neural network to reduce the dimension of the system from $N_{u}$ to a small latent state space. This is usually achieved by means of autoencoders [56, 75, 11, 53, 28] or other reduction techniques [66]. However, if $N_{u}$ is small the advantages of using a small latent space representation may not be counterbalanced by the increased cost of evaluating the neural network. In this work, we propose a solution to this latter problem by introducing an appropriate change of variables that significantly reduces the stiffness of the problem. Indeed, we will show that our method makes it possible to use explicit solvers to integrate neural ODEs representing stiff systems.
We report in Figure 2 a representation of the neural ODE. The hat symbol $\hat{\cdot}$ represents a suitable normalization of the data, see the next section for details. For the sake of simplicity, we will indicate in the captions the reference solution (the one that is comprised in the dataset) with the subscript “ref”, and the components of $\boldsymbol{u}$ as predicted by the neural ODE with $u_{i}$.
@@ 2.4.1 Training a Neural ODE
There are two main approaches to train a neural ODE. The first one is the classic supervised approach. Assuming that the value of $\boldsymbol{f}(t^{(n)},\boldsymbol{u}(t^{(n)};\boldsymbol{\mu}^{(j)});\boldsymbol{\mu}^{(j)})$ is known for each sample in the discrete manifold $\widetilde{\mathcal{M}}$, then it is possible to compute the MSE loss and minimize it through a gradient-based optimizer. However, this kind of data is not always available. Instead, the second approach – the one originally proposed for the training of neural ODEs – minimizes the error with respect to the timeseries data $\boldsymbol{u}(t^{(n)};\boldsymbol{\mu}^{(j)})$ for a fixed $j$. Namely, by using a differentiable ODE solver, we minimize the loss
@@
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{N_{\textnormal{train}}}\sum_{j=1}^{N_{\textnormal{train}}}\left\|{\boldsymbol{u}(t^{(N)};\boldsymbol{\mu}^{(j)})}-\left(\boldsymbol{u}(t^{(0)};\boldsymbol{\mu}^{(j)})+\int_{t^{(0)}}^{t^{(N)}}\mathcal{NN}(\tau,\boldsymbol{u}(\tau;\boldsymbol{\mu}^{(j)}),\boldsymbol{\mu}^{(j)};\boldsymbol{\theta})\mathrm{d}\tau\right)\right\|_{p},$ (6)
@@
where $\left\|\cdot\right\|_{p}$ indicates the discrete $p$-norm, the elements $\boldsymbol{u}(t^{(0)};\boldsymbol{\mu}^{(j)})$ and ${\boldsymbol{u}(t^{(N)};\boldsymbol{\mu}^{(j)})}$ belong to the solution manifold $\widetilde{\mathcal{M}}$, $\mathcal{NN}$ is the approximation of $\boldsymbol{f}$ made by a suitably trained neural network and the integral is usually computed with an explicit integration method like forward Euler or an explicit Runge-Kutta method.
@@
The advantage of the supervised approach is that it is less computationally demanding since it does not need to compute the gradient through the ODE solver using automatic differentiation. On the other hand, in the supervised framework, the loss is not always indicative of the accuracy of the predictions on the timeseries data $\boldsymbol{u}(t^{(n)};\boldsymbol{\mu}),n=0,...,N$. For this reason, we always employ as validation loss Eq. (6).
@@
Implementation details for Eq. (6) are important when it is employed to train the neural network. We resample the discrete timeseries to match the output of the time integrator, namely a Runge-Kutta method of the fourth order with a fixed timestep. To minimize the interpolation errors, we use a reference solution with a very fine time discretization. A uniform resampling allows for an efficient use of the partial results of the ODE solver for all the times smaller than the final one. Moreover, the initial and final integration times are not $(t^{(0)},t^{(N)})$ but are batched by uniform extraction between $[0,T]$ so that each time integration in the dataset is done with the same number of timesteps: this procedure enables vectorization of the computations, greatly reducing the computational cost. This optimization was originally introduced in torchdiffeq [20]. The number of timesteps used is called unroll length and is a hyperparameter that plays a key role in the optimization procedure. Finally, as commonly done in supervised learning, the computations are split into mini-batches to increase stochasticity. If not otherwise stated, when using Eq. (6) for validation or testing we employ as a time integrator an embedded Runge-Kutta method of order four with adaptive time-stepping. However, computations are significantly cheaper since we can turn off automatic differentiation.
@@
In our experiments, we employ a two-phase training. First, a standard supervised approach is used to “initialize” the neural network since it is cheap and enables a fast reduction of the loss. The derivative data is estimated with a filtering approach from the trajectory data (cf. Section 4). Then, we switch to the differentiable ODE solver approach to fine-tune the ROM. Our experiments show that in this case, it is beneficial to use an increasingly larger unroll length to ensure the robustness and accuracy of the model. The two-phase training is not strictly needed, but we have found it significantly reduces the time and effort required for hyperparameter tuning compared to using only the trajectory data. Further details will be given in Section 4 for each test case. Employing a two-phase training approach not only reduces the computational cost but also increases the quality of the final model. Indeed, from our experiments, we have observed that with the supervised learning approach the neural network is able to learn sharper gradients that may lead to unstable states. On the other hand, learning through the ODE solver produces stable results that might follow less precisely the reference solution. Figure 3 shows an example for two (underfitting) models. In Figure 4 we show the training and validation loss for a model trained in sequence with the supervised and ODE solver approaches.
@@
To accelerate the training it is key to apply suitable normalizations to the input $\boldsymbol{\mu}$, the output $\boldsymbol{f}$, and the timeseries data $\boldsymbol{u}$. The exact expressions for the normalizations depend on the test case considered and are detailed in Section 4. In some special cases, e.g., if $\boldsymbol{\mu}$ is high-dimensional, further treatment of $\boldsymbol{\mu}$ may be needed to accelerate the training. The insight behind the choice of the normalization is that we aim to have data that has a magnitude close to one, does not have outliers, and has either a uniform distribution or a normal distribution with a standard deviation of magnitude close to one. In our experiments, strictly monotone normalizations achieved better results. We designed the normalizations by manual inspection of the histograms of the data.
@@ 3 Addressing the stiffness issue
While many techniques have been developed to build reduced-order models of large systems stemming from semi-discretized PDEs, not many solutions are available for stiff ODEs. Indeed, it is a challenge to learn their complex dynamics, which often feature large gradients and widely different time scales. More rigorously, a commonly used formula to define the stiffness $S$ index of a problem is
$S=\frac{\mathfrak{R}(\lambda_{\max{}})}{\mathfrak{R}(\lambda_{\min{}})}T,$ (7)
@@
where $\lambda_{i}$ are the eigenvalues of the Jacobian of the system and $\mathfrak{R}$ indicates the real part. However, it is known that this definition is not always helpful. For instance, it does not apply to the classic Robertson chemical kinetics problem since it has two zeros eigenvalues [70]. A more pragmatic definition is given by Harier: “Stiff equations are problems for which explicit methods do not work[79]. That is, explicit numerical time integrators such as the embedded Runge-Kutta method of the fourth order require timesteps so small that implicit methods become cheaper. Stiffness is a particularly important issue when working with neural ODEs. Indeed, it is a well-known fact that neural ODE training may fail due to stiffness. For instance, [42] shows issues of this kind and proposes techniques to improve the training. However, it is still necessary to use a costly implicit solver in the online phase, since the learnt dynamic is stiff.
@@
Hence, to use the neural ODE as a ROM, we aim to reduce the stiffness of the system, allowing for the use of (cheap) explicit methods. To this end, we propose to change the dynamics of the problem by considering the change of variable in time induced by the adaptive time-stepping of the implicit solver used to solve the FOM. We remark that this change of variable is particularly effective in the case of stiffness characterized by a large $S$. However, in cases where stiffness is characterized by the presence of an oscillating fast solution, reparametrization techniques show several limitations [3, 67]. Our approach is not an exception, and this is an area for further improvement.
We suppose the existence of a change of variable $t_{s}=t_{s}(t)$ that makes the system less stiff. Namely, instead of learning Eq. (1), we learn the following system
$\begin{cases}\dot{\boldsymbol{u}_{s}}(t_{s};\boldsymbol{\mu})=\boldsymbol{f}_{s}(t_{s},\boldsymbol{u}_{s};\boldsymbol{\mu}),\qquad t_{s}\in(0,1]\\ \boldsymbol{u}_{s}(0;\boldsymbol{\mu})=\boldsymbol{u}_{s,0}(\boldsymbol{\mu}),\end{cases}$ (8)
where, $\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu})=\boldsymbol{u}(t(t_{s});\boldsymbol{\mu})$ and thus $\boldsymbol{f}_{s}$ is
$\boldsymbol{f}_{s}(t_{s},\boldsymbol{u}_{s};\boldsymbol{\mu})=\frac{\text{d}\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu})}{\text{d}t_{s}}=\frac{\text{d}\boldsymbol{u}(t(t_{s});\boldsymbol{\mu})}{\text{d}t}\frac{\text{d}t}{\text{d}t_{s}}=\boldsymbol{f}(t(t_{s};\boldsymbol{\mu}),\boldsymbol{u};\boldsymbol{\mu})\dot{t}(t_{s};\boldsymbol{\mu}).$
We aim to build the time mapping $t_{s}=t_{s}(t)$ in a data-driven manner, in particular we suppose that
$t_{s}(t^{(n)})=\frac{n}{N},$
@@
that is, the $\Delta t$ chosen at the $n$-th step by the implicit time stepper is proportional to the derivative of the function that we employ as a change of variable in time. We remark that $t_{s}$ is strictly increasing and thus invertible, with inverse $t=t(t_{s})$. We guess this change of variables induces less stiff dynamics since the adaptive time-stepping spreads the sharp peaks and quickly adapts to the large timescales. The insight behind this approach is that using an implicit solver to train the neural ODE on Eq. (1) has a cubic cost [42], instead, this change of variable allows the training via an explicit method, which is significantly cheaper. In other words, the adaptive timestepping is embedded into the neural ODE, avoiding the backpropagation step through the implicit solver. While the proposed approach is closely related to classical embedded methods, it reduces significantly the cost of training a neural ODE.
While it could be possible to learn $t(t_{s};\boldsymbol{\mu})$ and $\dot{t}(t_{s};\boldsymbol{\mu})$ by simple regression methods (such as monotonicity-preserving interpolation), this approach has two significant drawbacks for our use case: (i) it becomes unfeasible to use Eq. (6) to train the neural ODE; (ii) these interpolation schemes do not generalize well for $t>T$ and quickly become expensive as the dimension of $\boldsymbol{\mu}$ increases.
Thus, in this framework, we have two neural networks. The first, $\mathcal{NN}$ is used to approximate $\boldsymbol{f}_{s}$, which only requires pointwise estimates of these values. Further details for the training are given in Section 3.1. The second ingredient needed to accommodate for the time reparametrization is a way to compute $t(t_{s};\boldsymbol{\mu})$ in the online phase, to map the solution back to the original dynamics. Our experiments show that learning directly the map $t=t(t_{s};\boldsymbol{\mu})$ is a hard task for a neural network, probably due to the presence of abrupt changes in the function. Instead, we found that learning $\dot{t}(\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu}))$ is a stable approach that produces more accurate results. We note that, provided $t=t(t_{s};\boldsymbol{\mu})$ is continuously differentiable (or just absolutely continuous, if we only require the derivative almost everywhere), it can always be represented – and hence learned via the neural ODE approximation – as
$t(t_{s};\boldsymbol{\mu})=\int_{t_{s}^{(0)}}^{t_{s}}\dot{t}(\tau;\boldsymbol{\mu})\text{d}\tau\approx\int_{t_{s}^{(0)}}^{t_{s}}\mathcal{NN}_{t}(\tau,\boldsymbol{u}_{s}(\tau;\boldsymbol{\mu}),\boldsymbol{\mu})\text{d}\tau,$
by the fundamental theorem of calculus. In this case, $\boldsymbol{u}_{s}(\tau;\boldsymbol{\mu})$ is just an extra input used to accelerate the training. However, to improve the generalization properties and keep the system autonomous, we prefer to remove the explicit dependence on $t_{s}$. This is possible only if the trajectories of the ODE are an injective immersion in $\mathbb{R}^{N_{u}}$ or if the “velocity” $\dot{t}(t_{s};\boldsymbol{\mu})$ is uniquely well defined at every point of the trajectory. Since it is non-trivial to prove that this second property holds, we have only empirically tested the hypothesis by analyzing the data. Indeed, in practice, this approach works even if the trajectories satisfy the previous properties almost everywhere. Hence, in all our test cases, the time mapping is computed with the following path integral
$t(t_{s};\boldsymbol{\mu})\approx\int_{\boldsymbol{u}_{s}(t_{s}^{(0)})}^{\boldsymbol{u}_{s}(t_{s})}\mathcal{NN}_{t}(\boldsymbol{u}_{s};\boldsymbol{\mu})\text{d}\boldsymbol{u}_{s},$
@@
where the values of $\boldsymbol{u}_{s}$ are the ones obtained by solving Eq. (8) by time integration of the neural ODE $\mathcal{NN}$. Our experiments (see Section 4) demonstrate that the method generalizes well to parameters $\boldsymbol{\mu}$ not present in the training set. The evaluation of the time dynamics is cheap since it can be done in one batch using vectorization. Indeed, in the first step, we compute $\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu})$, then, we carry out the cumulative integration of $\dot{t}(\boldsymbol{u}_{s}(t_{s};\boldsymbol{\mu}))$ with Simpson’s rule. In our experiments, this step requires only a few milliseconds and its cost is negligible with respect to the cost of the integration of the neural ODE. The main advantage of this approach is that it is cheap compared to autoencoders, especially for small problems since the only overhead is the computation of $t=t(t_{s};\boldsymbol{\mu})$. However, it is also possible to combine the autoencoder with the described methodology: first, the autoencoder extracts a latent state representation, then, we could apply the time reparametrization in the latent space, reducing the stiffness of the latent dynamics.
Unlike traditional non-constant time stepping, our data-driven time reparametrization fundamentally alters the underlying time dynamics by introducing a new, normalized time variable $t_{s}$ based on the adaptive time-stepping behavior of the original system. While the results are in practice similar, this technique enables to predict the adaptive timestepping for a new model parameter $\boldsymbol{\mu}$ in a much cheaper way, without the necessity of evaluating the Jacobian or solving a linear system, like it is requested in an implicit ODE solver.